# ConfuGuard: A Low False-Positive System for Detecting Malicious and Stealthy Typosquatting Threats in Package Registries

### v2

@@ Abstract
Package confusion attacks such as typosquatting threaten software supply chains. Attackers make packages with names that syntactically or semantically resemble legitimate ones, tricking engineers into installing malware. While prior work has developed defenses against package confusions in some software package registries, notably NPM, PyPI, and RubyGems, gaps remain: high false-positive rates; generalization to more software package ecosystems; and insights from real-world deployment.
In this work, we introduce ConfuGuard, a solution designed to address the challenges posed by package confusion threats. We begin by presenting the first empirical analysis of benign signals derived from prior package confusion data, uncovering their threat patterns, engineering practices, and measurable attributes. We observed that 13.3% of real package confusion attacks are initially stealthy, so we take that into consideration and refined the definitions. Building on state-of-the-art approaches, we extend support from three to six software package registries, and leverage package metadata to distinguish benign packages. Our approach significantly reduces 64% false-positive (from 77% to 13%), with acceptable additional overhead to filter out benign packages by analyzing the package metadata. ConfuGuard is in production at our industry partner, whose analysts have already confirmed 301 packages detected by ConfuGuard as real attacks. We share lessons learned from production and provide insights to researchers.
@@ 1 Introduction
Figure 1: Threat model illustrating common active and stealth package confusion attacks. These malicious packages can mislead developers into installing them or being included as dependencies in downstream packages.
Software package registries (SPRs) are indispensable in modern development. They help engineers develop applications by integrating open-source packages, reducing costs and accelerating product cycles [45, 50, 65]. SPRs are used by industry and governments [10, 55], including in AI and safety-critical applications [26, 53]. The resulting software supply chain is an attractive target — if an adversary can cause an engineer to install a malicious package, the attacker may gain access to downstream users’ resources. One of their tactics is package confusion attacks [27, 39]: creating malicious packages with mis-leading names [22, 28, 42]. Figure 1 shows the attack model, covering naming tactics (lexical, syntactic, and semantic confusion) and attack timing (active vs. stealthy).
Researchers have proposed various methods to mitigate package confusion attacks. Early solutions detected packages with lexically similar names [63, 56, 48]. State-of-the-art approaches have begun exploring semantic similarity [39]. However, as criticized by Ohm et al. [41], prior work did not consider the false positive rate which is important for industry use [41]. We observe two other shortcomings: prior work focuses on limited SPRs, notably NPM and PyPI; and the literature lacks any experience reports from production deployment.
With our industry partner, we developed a novel system to detect package confusion attacks. To address the problem of high false positive rates, we measured the engineering practices of package confusion attackers, and propose signals, based on package metadata, that distinguish benign packages from genuine (though perhaps stealthy) threats. To generalize prior work to more SPRs, we analyzed package naming practices and confusion attacks in six SPRs, and identified novel aspects of these SPRs that expand the attack surface for confusion attacks. To make our approach production-grade, we develop embedding-based name comparisons and a novel nearest-neighbor search algorithm that provides greater efficiency and accuracy.
We integrated these techniques in the ConfuGuard system. We evaluated ConfuGuard on previous attack datasets as well as several months’ of production data. We report that ConfuGuard achieves a false positive rate of 13%, an improvement of 64% over the state of the art. On a modern server, ConfuGuard has average latency of 6.8 seconds per package. In production, ConfuGuard identified 301 confirmed package confusion attacks in its first three months.
@@
  • We generalize prior analyses to several new SPRs, noting the effect of package naming hierarchies on naming attack tactics.
  • We propose novel algorithms to improve attack detection accuracy (embedding) and efficiency (nearest neighbor search).
  • We open-source our production dataset, ConfuDB, which includes 1,561 production data triaged by security analysts.
@@
Significance: Package confusion attacks are common yet challenging to detect. This paper extends current knowledge by introducing a novel detection system, designed through empirical analysis of previous data and insights gained from ongoing deployment. ConfuGuard is used in production by our industry partner.
@@ 2 Background and Related Work
Here we discuss software package registries and naming conventions (section 2.1), package confusion attacks (section 2.2), and defenses (section 2.3).
@@ 2.1 SPRs and Package Naming
Much of modern software development depends on community ecosystems of libraries and frameworks that facilitate software reuse [65, 67]. We refer to these reusable software artifacts as software packages. These packages are distributed by software package registries (SPRs) [50, 47]. Table 1 shows the SPRs for six popular ecosystems comprising over 7 million software packages. Some SPRs restrict package publication to individual accounts, so the publisher’s identifier does not appear in the package name. Others allow grouping packages into namespaces or organizations, which are incorporated into the package name. We call the former “1-level” and the latter “hierarchical”.
SPRs typically act as infrastructure and leave it to the engineering community to establish norms for interaction. One aspect of their “hands-off” approach is that SPRs impose almost no constraints on package names — they define an allowed character set and forbid duplicate names [4, 7, 5, 6, 2, 3]. This flexibility introduces two related challenges for engineers. First, based on application needs (e.g., traditional software vs. machine learning), engineers in different SPRs have opted for different naming conventions, which can enhance readability, maintainability, and collaboration [17, 25], but which may also confuse newcomers. Second, as we detail next, SPRs’ permissive naming practices enables adversaries to choose package names to deliberately confuse engineers.
Table 1: Ecosystems examined in this study, highlighting their primary domains, naming conventions, and SOTA works on package confusion detections. Registry sizes are as of Jan. 2025. An example 1-level naming pattern is PyPI’s requests module. An example hierarchical name is Hugging Face’s google-bert/bert-base-uncased.
Registry (# pkgs) Domain Name Structure Prior Work
PyPI (619K) Python 1-level [63, 39]
RubyGems (212K) Ruby 1-level [39]
@@ 2.2 Package Confusion Attacks and Taxonomy
Package confusion is an attack in which adversaries seek to persuade engineers to install malicious packages [39]. [^1] Package confusion is the SPR-specific manifestation of the general naming confusion attack, which affects all IT endeavors including DNS (domain names) [19, 60, 21], , BNS (Blockchain names) [37], and container images (image names) [31, 8]. Attackers typically choose a package name similar to a legitimate one that meets the engineer’s needs [39]. Attacks may be multi-pronged, e.g., using advertising campaigns to increase their chances of fooling engineers [49, 24]. Package confusion attacks are an ongoing concern, with hundreds of packages detected by researchers in Socket [57], ReversingLabs [46], Orca [1], and others [13].
Table 2 summarizes the state-of-the-art taxonomy of package confusion techniques, based on Neupane et al. [39]. The variations in naming structure in different SPRs also affect the nature of package confusion by SPR. In SPRs with 1-level names, as shown in prior work, confusion attacks look like the top portion of the table. In hierarchical SPRs (this work), the attacker has a larger naming surface to exploit (bottom of table).
Table 2: Common package confusion techniques with examples. The top section presents some of the taxonomy proposed by Neupane et al. [39]. We added four patterns (bottom, cf. section 4.1), generalizing the prior taxonomy to attacks in other SPRs. The examples for those four patterns are confusion threats flagged by ConfuGuard.
@@ 2.3.1 Approaches
In the academic literature, the earliest name-oriented defenses were purely lexical: Taylor et al. [56] and Vu et al. [63] measured similarity using Levensthein distance [30] to identify PyPI packages with minor textual alterations from one another. To detect confusion attacks based on name semantics (e.g., bz2file vs. bzip), Neupane et al. [39] applied FastText embeddings [14] to enable an abstract comparison of package names.
In acknowledgment of package confusion attacks, industry has also offered some defenses. NPM and PyPI both use lexical distances to flag packages or otherwise reduce user errors [40, 44, 68]. Some other companies also offer relevant security tools. For example, Stacklok combines Levenshtein distance with repository and author activity metrics to assign a risk score to suspicious packages [52]. Microsoft’s OSSGadget generates lexical string permutations and verifies the existence of mutated package names on SPRs to detect potential package confusion attacks [36].
2.3.2 Knowledge gaps
Despite these advances, key challenges persist that hinder the accuracy and reliability of current package confusion detection approaches. The primary concern was articulated by Ohm et al. [41]: these techniques have high false-positive rates. Second, detailed empirical treatments have so far focused exclusively on package confusion attacks in NPM, PyPI, and RubyGems [56, 39, 63], leaving unknown the confusion attacks in SPRs with hierarchical naming structures. Finally, although summaries of package confusion defenses in production are reassuring, the lack of detailed production data makes their utility unclear.
@@ 3 Problem Statement
In this section, we give our definition of package confusion threats (section 3.1), our threat model (section 3.2), and our system requirements (section 3.3).
@@ 3.1 Refined Definition of Package Confusion
We refine the definition of package confusion threat by emphasizing the manifestation of malicious intent. Our refined definition is supported by our empirical analysis of real attacks in section 4.1.
@@
The mimicry may be lexical (i.e., typosquatting), syntactic, or semantic, but the intent is to confuse an engineer.
Because our definition considers threats, not just attacks, the current absence of malware does not imply future innocence. In contrast, prior work relies on malware detectors and only finds active attacks [56, 39]. Incorporating stealthy threats into this definition raises the possibility of high false positive rates. To address this, we analyzed false positives from prior package confusion data (section 4.2) to identify signals that distinguish benign packages from stealthy threats.
@@ 3.2 Threat Model
We describe the context and threats considered in this work.
Context Model: We focus on software package registries, including SPRs hosting traditional software packages and pre-trained AI models. The relevant properties are given in section 2.1.
Threat Model: We include some threats and exclude others.
  • In-scope: We consider attacks where packages with deceptively similar names to legitimate packages are published in the software package registry. As our definition permits, the attacker may initially publish non-malicious content (stealthy threat) and later introduce malware (active threat).
  • Out-of-scope: We exclude attacks using non-confusing package names, or that compromise existing legitimate packages. These attacks must be mitigated by other measures [42].
This threat model is stronger than those of existing package confusion detection techniques [63, 56, 39], because we permit attacks to start with stealthy (non-malicious) package content.
@@ 3.3 Requirements
A detection system must meet both security and operational goals. Working with our industry partner, we identified four requirements:
Req1: High precision and recall. A threat detection system must balance precision (low false positive rate) against recall (low false negative rate). False negatives enable attacks, but false positives lead to alert fatigue [11].
Req2: Efficient and Timely Detection. The system must scale to large registries — low-latency checks enabling real-time feedback.
Req3: Compatibility Across Ecosystems. The system must support many SPRs with their diverse naming schemes (section 2.1).
Req4: High Recall of Stealth Package Confusion Attacks. The system must identify both active and stealth package confusions.
@@ 4 Analysis of Confusion Attacks
To address the high false positive rates of previous detection systems for package confusion, Ohm et al. recently called for a deeper analysis of the available evidence [41]. We respond to their call in two ways: describing attackers’ practices (section 4.1), and reporting metadata features that distinguish true from false positives (section 4.2).
@@ 4.1 Attackers’ Practices
We examined first the role of stealth in confusion attacks, and then the attack variations in SPRs beyond NPM, PyPI, and RubyGems. Our findings informed requirements 1, 3, and 4 (section 3.3).
@@ 4.1.1 Stealthy Confusion Attacks
We analyzed the versions of the 240 confirmed NPM package confusion “true-positives”, defined as confusing packages that include malware. These data were collected and reported by Neupane et al. [39]. We analyzed the number of days these malicious packages were available before malware was injected. We estimated the injection time based on the last version updated before removal by the SPR.
Figure 2 illustrates our findings on the release time before malware injection. Most package confusion attacks injected malware on the first day of release, e.g., in the initial version. However, in 13.3% of the packages, attackers undertook stealth attacks, deploying malware after days or weeks. Attackers may use this strategy to grow a userbase before exploitation. Prior work on package confusion did not account for this behavior.

### v3

# ConfuGuard: Using Metadata to Detect Active and Stealthy Package Confusion Attacks Accurately and at Scale

@@ Abstract
Package confusion attacks such as typosquatting threaten software supply chains. Attackers make packages with names that syntactically or semantically resemble legitimate ones, tricking engineers into installing malware. While prior work has developed defenses against package confusions in some software package registries, notably NPM, PyPI, and RubyGems, gaps remain: high false-positive rates, generalization to more software package ecosystems, and insights from real-world deployment.
In this work, we introduce ConfuGuard, a state-of-art detector for package confusion threats. We begin by presenting the first empirical analysis of benign signals derived from prior package confusion data, uncovering their threat patterns, engineering practices, and measurable attributes. Advancing existing detectors, we leverage package metadata to distinguish benign packages, and extend support from three up to seven software package registries. Our approach significantly reduces false positive rates (from 80% to 28%), at the cost of an additional $14s$ average latency to filter out benign packages by analyzing the package metadata. ConfuGuard is used in production at our industry partner, whose analysts have already confirmed 630 real attacks detected by ConfuGuard.
@@ 1 Introduction
Figure 1: Threat model illustrating common active and stealth package confusion attacks. Attackers spoof a legitimate package such as bz2file with lexical typos (bz2fiel), syntactic namespace manipulations in scoped registries (@xx/bz2file), and semantic substitutions (bzip). Attackers release these look-alike packages as active attacks (malicious from the first version) or stealth attacks that are benign until a later update injects malware. Engineers may be tricked into depending on the impostor directly or as a transitive dependency.
Software package registries (SPRs) are indispensable in modern development. They help engineers develop applications by integrating open-source packages, reducing costs and accelerating product cycles [56, 63, 77]. SPRs are used by industry and governments [1, 67], including in AI and safety-critical applications [32, 65]. The resulting software supply chain is an attractive target — if an adversary can cause an engineer to install a malicious package, the attacker may gain access to downstream users’ resources. One of their tactics is package confusion attacks [35, 49]: creating malicious packages with misleading names [28, 37, 52]. Figure˜1 shows the attack model, covering naming tactics (lexical, syntactic, and semantic confusion) and attack timing (active vs. stealthy).
Researchers have proposed detectors to mitigate package confusion attacks. Early solutions detected packages with lexically similar names [75, 68, 61]. State-of-the-art approaches have begun exploring semantic similarity [49]. However, as criticized by Ohm et al. [51], prior work did not emphasize low false positive rates, which are needed for real-world use [51]. We note also that prior work focuses on limited SPRs, notably NPM and PyPI; and that the literature lacks experience reports from production deployment.
In this study we propose a novel system to detect package confusion attacks. To reduce false positive rates, we examine the practices of package confusion attackers and propose signals, based on package metadata, to distinguish benign packages from genuine (though perhaps stealthy) threats. To generalize prior work to more SPRs, we analyze naming practices and confusion attacks in seven SPRs, and identify novel aspects of these SPRs that expand the attack surface for package confusion. We also describe an embedding-based name comparison method and a novel nearest-neighbor search algorithm that improve the accuracy and efficiency of detection.
We integrated these techniques in the ConfuGuard system. We initially evaluated ConfuGuard on previous package confusion attack datasets, and iteratively enhanced it using recent real-world package confusion reports flagged and triaged by our industry partner’s security analysts. We report that ConfuGuard achieves a false positive rate of 28%, an improvement of 52% over the state of the art. ConfuGuard has average latency of 6.8 seconds per package on a modern server. Our industry partner has already used ConfuGuard in production to identify 630 confirmed package confusion attacks.
@@
  • We generalize prior analyses to several new SPRs, noting the effect of package naming hierarchies on confusion attack tactics.
  • We propose novel algorithms to improve the accuracy and efficiency of detecting package confusion attacks.
  • We open-source ConfuDB, our production dataset containing 2,361 package confusion threats triaged by security analysts.
@@
Significance: Package confusion attacks are common yet challenging to detect. This paper introduces ConfuGuard, a best-in-class detection system for package confusion detection, designed through analysis of previous attacks and insights from ongoing deployment. ConfuGuard is used in production by our industry partner.
@@ 2 Background and Related Work
Here, we provide background on SPRs and their package naming conventions (section˜2.1), review the taxonomy of package confusion attacks (section˜2.2), and outline existing defense mechanisms (section˜2.3).
@@ 2.1 SPRs and Package Naming
Much of modern software development depends on community ecosystems of libraries and frameworks that facilitate software reuse [77, 79]. We refer to these reusable software artifacts as software packages. These packages are distributed by software package registries (SPRs) [63, 59]. Table˜1 shows the SPRs for seven popular ecosystems comprising over 8 million software packages. Some SPRs restrict package publication to individual accounts, so the publisher’s identifier does not appear in the package name. Others allow grouping packages into namespaces or organizations, which are incorporated into the package name. We call the former package naming structure “flat” and the latter “hierarchical” [6].
Table 1: SPRs analyzed in this study, including their programming ecosystem, package naming structure, and presence in prior work on package confusion detection. Registry sizes are as of Jan. 2025. An example flat naming pattern is PyPI’s requests module. An example hierarchical name is Hugging Face’s google-bert/bert-base-uncased.
Registry (# pkgs) Ecosystem Name Structure Prior Work
PyPI (619K) Python Flat [75, 49]
RubyGems (212K) Ruby Flat [49]
@@
NuGet (430K) .NET Both N/A
SPRs typically act as infrastructure and leave it to the engineering community to establish norms for interaction. One aspect of their “hands-off” approach is that SPRs impose almost no constraints on package names — they define an allowed character set and forbid duplicate names [16, 17, 3, 13, 23, 42]. This flexibility introduces two related challenges for engineers. First, based on the application domain (e.g., traditional software libraries vs. machine learning model packages), engineers across different SPRs have adopted different naming conventions, which can enhance readability, maintainability, and collaboration [14, 31], but which may also confuse newcomers. Second, as we detail next, SPRs’ permissive naming practices enable adversaries to choose package names to deliberately confuse engineers.
@@ 2.2 Package Confusion Attacks and Taxonomy
Package confusion is an attack in which adversaries seek to persuade engineers to install malicious packages [49]. [^1] Package confusion is the SPR-specific manifestation of the general naming confusion attack, which affects all IT endeavors including DNS (domain names) [18, 72, 24], BNS (Blockchain names) [47], and container images (image names) [40, 5]. Attackers typically choose a package name similar to a legitimate one that meets the engineer’s needs [49]. Attacks may be multi-pronged, e.g., using advertising campaigns to increase their chances of fooling engineers [62, 30]. Package confusion attacks are an ongoing concern, with hundreds of packages detected by researchers in Socket [69], ReversingLabs [58], Orca [12], and others [7].
Table˜2 summarizes the state-of-the-art taxonomy of package confusion techniques, based on Neupane et al. [49]. The variations in naming structure in different SPRs also affect the nature of package confusion by SPR. In SPRs with a flat naming structure, confusion attacks look like the top portion of the table. In hierarchical SPRs, attackers have a larger naming surface to exploit, as illustrated in the bottom portion of the table.
Table 2: Common package confusion techniques with examples. The top section presents some of the taxonomy proposed by Neupane et al. [49]. We extend this taxonomy by adding three new patterns (bottom; see section˜4.1) that generalize confusion strategies observed in additional SPRs. The examples for those new patterns are drawn from package confusion threats flagged by ConfuGuard.
@@ 2.3.1 Approaches
In the academic literature, the earliest name-oriented defenses were purely lexical: Taylor et al. [68] and Vu et al. [75] measured similarity using Levenshtein distance [39] to identify PyPI packages with minor textual alterations from one another. To detect confusion attacks based on name semantics (e.g., bz2file vs. bzip), Neupane et al. [49] applied FastText embeddings [9] to enable an abstract comparison of package names.
Recognizing the threat of package confusion attacks, industry has also offered some defenses. NPM and PyPI both use lexical distances to flag packages or otherwise reduce user errors [50, 55, 80]. Some other companies also offer relevant security tools. For example, Stacklok combines Levenshtein distance with repository and author activity metrics to assign a risk score to suspicious packages [64]. Microsoft’s OSSGadget generates lexical string permutations and verifies the existence of mutated package names on SPRs to detect potential package confusion attacks [46].
Except for Stacklok’s, these defenses rely solely on the characters in the package name, whether through edit-distance heuristics or semantic similarity. We refer to these lexical or semantic name features collectively as “string-only signals”. These defenses ignore auxiliary metadata such as popularity, maintainers, and READMEs.
2.4 Gaps in Current Approaches
Despite the proposed approaches for detecting package confusion attacks, both theoretical and practical gaps remain.
2.4.1 Theoretical Gaps
Current detectors primarily depend on string-only signals, disregarding contextual metadata. This narrow view is likely to inflate false positive rates when legitimate projects share similar names, while allowing stealthy threats to go undetected when attackers use benign-looking package names combined with adversarially crafted metadata (e.g., package descriptions, READMEs) to deceive engineers. Furthermore, prior work’s definition of package confusion requires packages to contain active malware [49], failing to account for stealthy threats that initially appear benign but may become malicious later. This oversight highlights a gap in understanding the full threat landscape.
Additionally, detailed empirical treatments have so far focused exclusively on package confusion attacks in NPM, PyPI, and RubyGems [68, 49, 75], omitting package confusion attacks in SPRs with hierarchical naming structures such as Maven, Golang, and Hugging Face. This limited scope means that existing detection algorithms and similarity metrics may not generalize to SPRs with hierarchical names, which offer attackers additional attack surfaces through namespace manipulation and compound squatting techniques.
2.4.2 Practical Gaps
The primary concern, as articulated by Ohm et al. [51], is that current package confusion detectors have high false-positive rates. Prior studies have shown that engineers consider monitoring systems with excessive false positives to be less useful than conservative ones [20]. In the context of package confusion detection, false positives cause several practical issues: they lead to alert fatigue among security analysts, reduce the likelihood that real threats will be investigated; erode trust in the detection system, causing organizations to disable or ignore alerts; and harm legitimate package developers’ reputations [21].
The engineering community acknowledges the severity of false positives. For example, developers of Microsoft’s OSSGadget typosquatting detection tool explicitly seek ways to reduce false positives [60]. Additionally, while industry reports mention package confusion defenses being used in production (e.g., NPM and PyPI’s lexical distance checks, Stacklok’s risk scoring), these reports lack concrete deployment details such as false positive rates, detection accuracy, number of threats caught, or operational challenges. Without concrete performance metrics and deployment experiences, it remains unclear how effective these systems are in practice.
@@ 3 Problem Statement
In this section, we present an improved definition of package confusion threats (section˜3.1), outline our threat model (section˜3.2), and specify the system requirements necessary to effectively detect package confusion attacks (section˜3.3).
@@ 3.1 Refined Definition of Package Confusion
We refine the definition of package confusion threat by emphasizing the manifestation of malicious intent. Our refined definition is supported by our analysis of package confusion attacks (section˜4.1).
@@
As depicted in Figure˜1, the mimicry may be lexical (i.e., typosquatting), syntactic, or semantic, but the intent is to confuse an engineer.
By framing our definition around package confusion threats rather than just attacks, we admit the possibility that packages may initially appear benign but harbor malicious intent. In contrast, prior work relies on malware detectors and only finds actively malicious packages [68, 49]. Expanding the package confusion detection scope from active attacks to stealthy threats enlarges the candidate set of similar package pairs and, if done naively, can inflate false positives. We mitigate this challenge through our empirical analysis of false positive patterns in existing package confusion datasets (section˜4.2), which reveals distinguishing signals between genuinely benign packages and stealthy threats awaiting activation.
@@ 3.2 Threat Model
We define the scope and boundaries of our threat model to clarify which package confusion attacks are addressed. Our model encompasses both the context in which these attacks occur and the specific threats we aim to detect.
Context Model: We focus on the SPRs mentioned in section˜2.1, including both traditional SPRs hosting software packages and emerging platforms for pre-trained AI models.
Threat Model: Our threat model distinguishes between package confusion attacks we address and those outside our scope:
  • In-scope: We consider attacks where adversaries publish packages with names deceptively similar to legitimate packages within SPRs. Consistent with our refined definition, attackers may initially upload benign content (stealthy threats) and later inject malicious code (active threats).
  • Out-of-scope: We exclude attacks that either use non-confusing package names or compromise existing legitimate packages through other means. Such attacks require different mitigation strategies [52].
This threat model is stronger than those of existing package confusion detection approaches [75, 68, 49] because we account for stealthy attacks that begin with non-malicious package content.
@@ 3.3 Requirements
A system for detecting package confusion threats must meet both security and operational goals. Working with our industry partner, we identified four requirements:
Req1: High Precision and Recall. The system must balance precision (low false positive rate) against recall (low false negative rate). False negatives leave deceptive packages unflagged, but false positives lead to alert fatigue [2].
Req2: Efficient and Timely Detection. The system must scale to large SPRs while making low-latency checks for real-time feedback.
Req3: Compatibility Across Ecosystems. The system must support many SPRs with their diverse naming structures.
Req4: High Recall of Stealth Package Confusion Attacks. The system must identify both active and stealthy package confusions.
@@ 4 Analysis of Confusion Attacks
To address the high false positive rates of previous package confusion detection systems, Ohm et al. recently called for a deeper analysis of the available evidence [51]. We respond to their call in two ways: describing attackers’ practices (section˜4.1), and reporting metadata features that distinguish true from false positives (section˜4.2).
@@ 4.1 Attackers’ Practices
We examined first the role of stealth in confusion attacks, and then the attack variations in SPRs beyond NPM, PyPI, and RubyGems. Our findings informed Req1,3,4.
@@ 4.1.1 Stealthy Confusion Attacks
We analyzed 240 confirmed NPM package confusion “true positives” (i.e., confusing packages that include malware), based on data provided by Neupane et al. [49]. We analyzed the number of days these malicious packages were available before malware was injected. We estimated the injection time based on the last version updated before removal by the SPR.
Figure˜2 illustrates our findings on the release time before malware injection. Most package confusion attacks injected malware on the first day of release (i.e., in the initial version). However, in 13.3% of the packages, attackers undertook stealth attacks, deploying malware after days or weeks. Attackers may use this strategy to grow a user base before exploitation. Prior work on package confusion detection did not account for this behavior.